Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below:

  

Sales (12,800 units × $20 per unit) $ 256,000
Variable expenses 153,600
Contribution margin 102,400
Fixed expenses 114,400
Net operating loss $ (12,000 )

2. The president believes that a $6,400 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $83,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income?

3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $37,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)?
Answer #1
Req 2 :
Contribution margin ratio = Contribution margin / Sales = 102400 / 256000 40%
Increase in Contribution margin ( Increase in sales * Contribution margin ratio = 83000 * 40% ) 33200
(-) Increase in advertising budget 6400
Increase (decrease) in net operating income 26800
Increase by 26800
Req 3 :
Revised selling price = Current selling price * ( 1 - % reduction ) = 20 * ( 1 - 10% ) 18
Current unit variable cost = Variable expenses / Units sold = 153600 / 12800 12
Revised fixed costs = Current fixed costs + Increase in advertising expense = 114400 + 37000 151400
Revised sales units = Current sales units * 2 = 12800 * 2 25600
Calculation of revised net operating income (loss) :
Sales ( 25600 * 18 ) 460800
(-) Variable expenses ( 25600 * 12 ) 307200
Contribution margin 153600
(-) Fixed expenses 151400
Revised net operating income (loss) 2200